OpenStudy (anonymous):

A ship is anchored off a long straight shoreline that runs north and south. From two observation points 17 miles apart on shore, the bearings of the ship are N 32° E and S 53° E. What is the shortest distance from the ship to the shore? (Round your answer to the nearest tenth.) Please explain to me how to solve this problem. Thank you!

OpenStudy (anonymous):

From the attached figure. 1. The distance between the two observers A and B is 17. 2. shortest distance to shore is Z 3. x + y = 17 => x = 17 - y 4. z/y = tan(32) => z = y*tan(32) 5. z/x = tan(53) => z = x*tan(53) combine step 4 and 5 to get: 6. y*tan(32) = x*tan(53) from step 3, x = 17 - y combine step 6 and step 3 to get: 7. y*tan(32) = (17 - y)*tan(53) => y= [17 tan(53)]/[tan(32) + tan(53)] combine step 4 and step 7 to get: 8. Z={17*tan(53) *tan(32)}/[tan(32) + tan[53)] = 7.22209205254024310884

OpenStudy (anonymous):

Oh! That makes a lot of sense. Thank you so much for taking the time out to help me. I appreciate it!